Topic: Optimization


1. A close right circular cylinder is to be constructed to hold a 1 liter oil can shape.


What dimensions will minimize the amount of material, assuming that the


thickness of the material is uniform?



2. Find two positive numbers whose sum is 9 and whose product is a maximum.

1
Expert's answer
2022-05-31T10:10:55-0400

1. Let "r=" the radius of the base of the cylinder, let "h=" the height of the cylinder. Then


"V=\pi r^2h=1000 {cm}^3"

"h=\dfrac{1000}{\pi r^2}"

The total surface area of cylinder is


"A=2(\pi r^2)+2\pi rh"

Substitute


"A=A(r)=2\pi r^2+2\pi r(\dfrac{1000}{\pi r^2})"

"=2\pi r^2+\dfrac{2000}{ r}"

Differentiate with respect to "r"


"A'(r)=4\pi r-\dfrac{2000}{ r^2}"

Find the critical number(s)


"A'(r)=0=>4\pi r-\dfrac{2000}{ r^2}=0"

"r=\sqrt[3]{\dfrac{2000}{4\pi}}"

"r=5\sqrt[3]{\dfrac{4}{\pi}}"

If "0<r<5\sqrt[3]{\dfrac{4}{\pi}}, A'(r)<0, A(r)" decreases.

If "r>5\sqrt[3]{\dfrac{4}{\pi}}, A'(r)>0, A(r)" increases.

The function "A(r)" has the local minimum at "r=5\sqrt[3]{\dfrac{4}{\pi}}."

Since the function "A(r)" has the only extremum, then the function "A(r)" has the absolute minimum at "r=5\sqrt[3]{\dfrac{4}{\pi}}."


"h=\dfrac{1000}{\pi (5\sqrt[3]{\dfrac{4}{\pi}})^2}=10\sqrt[3]{\dfrac{4}{\pi}}"

"r=5\sqrt[3]{\dfrac{4}{\pi}}\ cm, h=10\sqrt[3]{\dfrac{4}{\pi}}\ cm"


2.

Let "x=" the first number. Then the second number will be "9-x."

Since two numbers are positive, then "0<x<9."

The product of two numbers is "f(x)=x(9-x)."


"f(x)=-x^2+9x"

This quadratic function has the absolute maximum at


"x=-\dfrac{9}{2(-1)}=\dfrac{9}{2}"

The first number is "\dfrac{9}{2}." The second number is "\dfrac{9}{2}."